Sandra would like to organize BAL as either an LLC (taxed as a sole proprietorship) or a C corporation. In either form, the entity is expected to generate an 8 percent annual before-tax return on a $500,000 investment. Sandra’s marginal income tax rate is 37 percent and her tax rate on dividends and capital gains is 23.8 percent (including the 3.8 percent net investment income tax). If Sandra organizes BAL as an LLC, she will be required to pay an additional 2.9 percent for self-employment tax and an additional .9 percent for the additional Medicare tax. BAL’s income is not qualified business income (QBI) so Sandra is not allowed to claim the QBI deduction. Assume that BAL will distribute all of its after-tax earnings every year as a dividend if it is formed as a C corporation.

a. How much cash after taxes would Sandra receive from her investment in the first year if BAL is organized as either an LLC or a C corporation?
Answer #1
LLC Description C.corp Description
1 Pretax Earnings 25000 (500000*8%) 25000 (500000*8%)
2 Entity level tax rate 0% 21%
3 Entity level tax 0 5250 (1)*(2)
4 Earnings after entity level tax 25000 (1)-(3) 19750 (1)-(3)
5 Deduction for 50% of SE tax (338) (1)*0.9325*0.029*0.5 NA
6 Net Income taxable to owner 24662 (4)+(5) 19750 (4) distributed as dividend
7 Income tax paid by owner 9125 (6)*0.37 3950 (6)*0.20
8 Self-Employment tax 670 (1)*0.9235*0.029
9 Additional Medical tax 208 (1)*0.9235*0.009 751 (6)*0.038 net investment income tax
10 Owner level tax 10003 (7)+(8)+(9) 4701 (7)+(9)
11 After tax cash flow 14997 (1)-(10) 5250 (6)-(10)
Overall tax rate 40.012% (10)/(1) 39.80% [(3)+(10)]/1

Part A:

After tax cash flow
LLC 14997
C corporation 5250

Part B:

Overall tax rate = 10003/25000=40.012%    (LLC)

=(4701+5250)/25000=39.80% ( C Corporation)